OpenStudy (anonymous):

One side of a parallelogram has endpoints (3, 3) and (1, 7). What are the endpoints for the side opposite? (8, 1) and (6, 7) (6, 1) and (2, 3) (6, 1) and (8, 5) (8, 1) and (6, 5)

OpenStudy (yttrium):

Do you remember your formula for slope using two point form?

OpenStudy (anonymous):

y – y1 = m(x – x1)

OpenStudy (yttrium):

No, it's not that.

OpenStudy (anonymous):

Again, I have plotted the graph for you. You should be able to complete your answer, **The three will look like a spec on your screen, but it is there. ;)** \(\color{purple}{Chance}\)

OpenStudy (anonymous):

still confused....

OpenStudy (yttrium):

let me tell you the formula?

OpenStudy (anonymous):

what's the formula?

OpenStudy (anonymous):

It is actually quite simple, really. I can plot each answer choice conjoing with the information given if you like. That sould make it easier for you to pick an answer. Would you like me to?

OpenStudy (anonymous):

Yes...plotting it will make it easier and ill figure it out myself

OpenStudy (yttrium):

m = (Y2 - Y1)/(X2 - X1) find first the slope of the line, and then use substitution for each point to find the new coordinates. Since parallelogram has parallel sides, you will use the same slope. Got it?

OpenStudy (anonymous):

Data given plotted with answer choice A.

OpenStudy (anonymous):

(8,1) (6,7) is the answer ?

OpenStudy (anonymous):

I'm not sure. I am just plotting them.

OpenStudy (anonymous):

Data given plotted with answer choice B.

OpenStudy (anonymous):

Data given plotted with answer choice C.

OpenStudy (anonymous):

Data given plotted with answer choice D.

OpenStudy (anonymous):

All of your choices have been plotted. It should be easy for you to figure out the answer.

OpenStudy (anonymous):

i have that same question:(